Simplify (4-2i)(5+i)


1
Expert's answer
2021-02-23T11:16:32-0500

(4-2i)(5+i)=-2(i-2)(i+5)=-2(i²-2i+5i-10)=-2i²+4i-10i+20=-2i²-6i+20